OpenStudy (anonymous):

Suppose a computer virus begins by infecting 8 computers in the first hour after it is released. Each hour after that, each newly infected computer causes 8 more computers to become infected. The function y = 8x models this situation. Make a table with integer values of x from 1 to 4 in the space below.

OpenStudy (anonymous):

\[y=8^x\] i assume

OpenStudy (anonymous):

f(x) x 8 1 16 2 24 3 32 4 like this?

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\(x \) \( f(x)\) \(1\) \(8^1=8\) \(2\) \(8^2=64\) \(3\) \(8^3=512\) \(4\) \(8^4=4096\)

OpenStudy (anonymous):

is that the way you write the table?

OpenStudy (anonymous):

that is one way, yes

OpenStudy (anonymous):

put the \(x\) first and the \(y\) second